OpenStudy (anonymous):

Which of the following is the correct equation for this function? it shows a graph with a parabola facing upwards with a vertex of (-3,-2) and the x intercept of (-2,0) and (-4,0) A. y=x^2+3x-2 B. y=2=-2(x+3)^2 C. y=(x+4)(x+2) D. y+2=2(x+3)^2

OpenStudy (anonymous):

option B looks invalid - can you check it please?

OpenStudy (anonymous):

ph my bad its supposed to be y+2=-2(x+3)^2

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thanks just wondering how did you get that answer so i know how to do it in future

OpenStudy (anonymous):

Substitute x = -2 in to the equation and you get y = 0. Substitute x = -4 in to the equation and you get y = 0 So it has x intercept of (-2,0) and (-4,0)

OpenStudy (anonymous):

But this is not the only one of the options for which this is true so we need to check the vertex as well

OpenStudy (anonymous):

so you need to simplify the function, expanding out the brackets, and get it in the form of y = ...... which is y= 2x^2+12x+16

OpenStudy (anonymous):

differentiate this to get dy/dx = 4x + 12 At the vertex, dy/dx = 0 so 4x + 12 = 0 so 4x = -12 so x = -3

OpenStudy (anonymous):

Sub this in to y= 2x^2+12x+16 to get y = -2 so the vertex is -3, -2) D fulfils all conditions specified so its the answer!

OpenStudy (anonymous):

cool thanks alot! :)

OpenStudy (anonymous):

you are most welcome :-) TheMathsTutor.Com